# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ 5 Numerical Experiments
We perform numerical experiments to evaluate MIN-UCB. We compare MIN-UCB with three related benchmarks. The first policy is vanilla UCB (Auer et al., 2002) that ignores the offline data, which we call “PURE-UCB”. The second policy chooses an arm with the largest $\text{UCB}^{\text{S}}_{t}(a)$ defined in (6), which we call “UCBS”. The third policy is HUCB1 (Shivaswamy & Joachims, 2012), which we call ”HUCB”.
In all experiments, we simulate our algorithm MIN-UCB and benchmarks on several instances, with fixed $K=4$. Each arms’s offline and online reward follows standard Gaussian distribution with mean $\mu^{\text{off}}(a)$, $\mu^{\text{on}}(a)$, respectively. In each specific instance, we run each algorithms 100 times and report their average and standard error. For simplicity, we assume $V(a)=V$ for any sub-optimal arm $a$ and $T_{\text{S}}(a)=T_{\text{S}}$ for any arm $a$. The detailed setup is provided in Appendix D.
Figure 1(a) plots the regret on single instance under different valid bias bounds $V$, with $T=T_{\text{S}}=10000$ in each trial. Specifically, for $V=0.2,0.6$, we illustrate the regret across different $T$ in Figures 1(b), 1(c). These results demonstrate the benefit of our algorithm. When $V$ is small, indicating that DM has a precise estimation on $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$, our MIN-UCB and UCBS that effectively leverage the information from $V$ and offline dataset perform well. In contrary, when $V$ is large, implying a loose estimation of $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$, the offline dataset becomes challenging to utilize effectively. In such cases, MIN-UCB’s performance aligns with PURE-UCB, and remains unaffected by looseness of $V$ that misleads the DM to believe $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ to be much larger than the actual gap. The stable performance of MIN-UCB is in contrast with the performance of UCBS and HUCB. Thus, the robustness of our algorithm is validated through these numerical results.
Figures 2(a) and 2(b) show the regret on different $T_{\text{S}}$, where in the first group $V=0.1$ (small $V$) and the second group $V=0.3$ (large $V$). Both figures demonstrate our MIN-UCB consistently outperforms alternative methods, irrespective of $T_{\text{S}}$ and $V$. A noteworthy observation is that is that when the offline data is easy to leverage (small $V$), a larger offline dataset (increased $T_{\text{S}}$) results in larger improvement of our algorithm over vanilla UCB. This observation aligns with intuition, as a larger $T_{\text{S}}$ coupled with a smaller $V$ indicates more informative insights into the online distribution. Conversely, when $V$ is large and offline data is hard to utilize, increasing the size of the offline dataset ($T_{\text{S}}$) does not significantly impact the performance of MIN-UCB. This is because our algorithm can adapt to this environment and consistently align with vanilla UCB, thereby mitigating the influence of the offline data on performance.
6 Conclusion
We explore a multi-armed bandit problem with additional offline data. Our approach on how to adapt to the possibly biased offline data is novel in the literature, leading to tight and optimal regret bounds. There are some interesting future directions. For example, we can further investigate how to apply it in other online learning models, such as linear bandits and online Markov decision processes. It is also intriguing to study how other forms of metrics that can measure the distribution drift can affect and facilitate the algorithm design.
Impact Statements
This paper presents work whose goal is to advance the field of Online Learning. We primarily focus on theoretical research. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A More Detailed Comparisons with Existing Works
A.1 Comparing with (Zhang et al., 2019)
(Zhang et al., 2019) consider a stochastic contextual $K$-armed bandit model with possibly biased offline data, which generalizes our setting of stochastic $K$-armed bandits. On one hand, their proposed algorithm ARROW-CB does not require any knowledge on the discrepancy between the offline and online data, while our proposed algorithm requires knowing an upper bound $V(a)$ to $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ for each $a$. On the other hand, in the case of stochastic $K$-armed bandits, the regret bound of ARROW-CB (see Theorem 1 in (Zhang et al., 2019)) is at least the regret bound of the explore-then-commit algorithm (see Chapter 6 in (Lattimore & Szepesvári, 2020)) that ignores all offline data, which in particular does not offer improved regret bound compared to existing baselines that do not utilize offline data.
@@
The function $\text{disc}(\lambda,I)$ (see Equation (4) in (Zhang et al., 2019)) is a discrepancy measure on the distance between the offline and online models, with the property that $\text{disc}(\lambda,I)\geq 0$ for all $\lambda,I$.
@@
Finally, one additional distinction is that (Zhang et al., 2019) requires $T_{\text{S}}(1)=\ldots=T_{\text{S}}(K)$, while we allow the offline sample sizes $T_{\text{S}}(1),\ldots,T_{\text{S}}(K)$ to be arbitrary non-negative integers.
@@ A.2 Comparing wtih Online Learning with Advice
The frameworks of (Rakhlin & Sridharan, 2013a; Steinhardt & Liang, 2014; Wei & Luo, 2018; Wei et al., 2020a) utilize a commonly defined quantity ${\cal E}$ to quantify the error in the predictions, which are provided in an online manner. We focus our discussion on comparison with (Wei & Luo, 2018). The framework of (Wei & Luo, 2018) provides the following results for adversarial multi-armed bandits with side information. Consider a $K$-armed bandit model, where an arm $a\in{\cal A}$ is associated with loss $\ell_{t}(a)\in[-1,1]$ at time $t$ for $t\in\{1,\ldots,T\}$, and the loss vectors $\{\ell_{1},\ldots,\ell_{T}\}$ over the horizon of $T$ time steps are generated by an oblivious adversary. Before choosing an arm $a_{t}\in{\cal A}$ at time $t$, the agent is endowed with a prediction $m_{t}\in[-1,1]^{K}$, where $m_{t}(a)$ is a prediction on $\ell_{t}(a)$. They design a version of optimistic online mirror design algorithm that incoporates the predictions $m_{1},\ldots,m_{T}$, and achieves an expected regret bound of
@@
We argue that their framework can only guarantees a $O(\sqrt{KT\log(T)})$ in terms of expected regret bound in our setting, since the prediction error term ${\cal E}$ depends on the realized losses rather than their means. Indeed, In our stochastic setting (translated to a loss minimization instead of reward maximization model), the loss vectors are $\ell_{1},\ldots,\ell_{T}$ are independent and identically distributed according to a common (but latent) distribution ${\cal D}$. Consider the case when the loss of arm $a$ is distributed according to the Bernoulli distribution with mean $\mu(a)\in[1/4,3/4]$ for each $a\in{\cal A}$. Even in the case when we have the prediction $m_{t}(a)=\mu(a)$ for each $t$, namely the predictions reveal the true mean, we still have $(\ell_{t}(a)-\mu(a))^{2}\geq 1/16$ for each arm with certainty, leading to ${\cal E}\geq T/16$.
The framework of (Wei & Luo, 2018) can only gauranttee a regret bound of $o(\sqrt{KT})$ when ${\cal E}=o(T)$. In our setting, it requires that $\|\ell_{t}-m_{t}\|_{\infty}=o(1)$, meaning that the prediction $m_{t}$ has to be correlated with, thus contains information of, the realized loss $\ell_{t}$, for $\Omega(T)$ many time rounds $t$. Different from (Wei & Luo, 2018), we achieve a regret bound of $o(\sqrt{KT})$ in the presense of accurate offline data, instead of receiving hints about the realized online rewards. A similar conclusion (with a worse regret bound than (18)) holds when we compare our results with (Wei et al., 2020a), which presents regret bounds on adversarial contextual bandits with predictions, since their regret bounds are also defined in terms of ${\cal E}$ mentioned above.
A.3 Comparing with (Chen et al., 2022)
(Chen et al., 2022) explores a Finite-Horizon Reinforcement Learning Model incorporating possibly biased historical data, another generalization of our stochastic multi-armed bandit model. Similar to our work, their approach ”Data-pooling Perturbed LSVI” (Algorithm 2) requires an upper bound on the discrepancy between offline and online data, defined as $\Delta$ in their Assumption 1. They propose an cross-validation method and illustrate through a case study that a substantial range of $\Delta$ values produce similar performance for their algorithm. However, their approach cannot facilitate the design of an algorithm that achieve our improved regret bounds compared to existing benchmarks that neglects the offline data.
@@
However, (19) does not offer an explicit closed-form bound, and they do not provide any regret lower bound.More importantly, their implicit bound solely depends on $T_{\text{S}}^{\min}=\min_{h,s,a}T_{\text{S}}(h,s,a)$, which is equivalent to $\min_{a}T_{\text{S}}(a)$ in our model. In contrast, both our instance dependent bound (Theorem 4.2) and instance independent bound (Theorem 4.9) demonstrate how the difference in $T_{S}(a)$ among different arms affects the regret. Thus, their approach cannot deliver the tight regret bounds on and insights into our model.
Appendix B Proofs for Regret Upper Bounds
B.1 Proof for Lemma 4.1
@@ Appendix B Proofs for Regret Upper Bounds
Proposition B.1 (Chernoff Inequality).
@@
Proof of Lemma 4.1.
To prove the Lemma, it suffices to prove that $\Pr[{\cal E}^{\text{S}}_{t}(a)]\geq 1-\delta_{t}$. Indeed, the inequality implies $\Pr({\cal E}_{t}(a))\geq 1-\delta_{t}$ by setting $T_{\text{S}}(a)=0$, and then a union bound on the failure probabilities of ${\cal E}^{\text{S}}_{t}(a),{\cal E}_{t}(a)$ for $a\in{\cal A}$ establishes the Lemma. Now, we have
@@
$\geq$ $2\displaystyle 1-\delta_{t}/$ 2 . (24)
The justifications of (23, 24) are the same as (21, 22) respectively. Altogether, the Lemma is proved.
B.2 Proof of instance dependent regret upper bound
B.2.1 Proof of Lemma 4.8
We consider two main cases depending on $\omega(a)$ and $\Delta(a)$:
Case 1: $\omega(a)<\Delta(a)$. The case condition implies that
$\max\left\{1-\frac{\omega(a)}{\Delta(a)},0\right\}^{2}=\left(1-\frac{\omega(a)}{\Delta(a)}\right)^{2}>0,\quad\frac{\omega(a)}{\Delta(a)}\in[0,1).$
To proceed. we consider two sub-cases:
Sub-case 1a: $T_{\text{S}}(a)\cdot(1-\frac{\omega(a)}{\Delta(a)})^{2}\geq\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$. Then
$\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}\leq\sqrt{\frac{2\log(4Kt^{4})}{T_{\text{S}}(a)}}\leq\left(1-\frac{\omega(a)}{\Delta(a)}\right)\sqrt{\frac{2\log(4Kt^{4})}{\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}}}=\left(1-\frac{\omega(a)}{\Delta(a)}\right)\cdot\frac{\Delta(a)}{\sqrt{8}}.$ (25)
Consequently, conditioned on event ${\cal E}_{t}$,
$\text{UCB}^{\text{S}}_{t}(a)$ $\leq\mu^{\text{(on)}}(a)+2\cdot\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)+T_{\text{S}}(a)}}+\frac{T_{\text{S}}(a)}{T_{\text{S}}(a)+N_{t}(a)}\cdot\omega(a)$ (26)
$<\mu^{(\text{on})}_{*}\leq\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$ (27)
The strict inequality in step (27) is by the sub-case condition that $01-\frac{\omega(a)}{\Delta(a)}>$ 0. The less-than-equal steps in Lines (26, 27) is by conditioning on the event ${\cal E}_{t}$. Altogether, $A_{t}\neq a$ with certainty.
Sub-case 1b: $T_{\text{S}}(a)\cdot(1-\frac{\omega(a)}{\Delta(a)})^{2}<\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}$. In this case, we clearly have
$N_{t}(a)>32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}-T_{\text{S}}(a)\cdot\max\left\{1-\frac{\omega(a)}{\Delta(a)},0\right\}^{2}\geq\frac{16\log(4Kt^{4})}{\Delta(a)^{2}}.$
Consequently, we have
leading to
$\text{UCB}_{t}(a)\leq\mu^{\text{(on)}}(a)+2\cdot\frac{\Delta(a)}{\sqrt{8}}<\mu^{(\text{on})}_{*}\leq\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$
thus $A_{t}\neq a$ with certainty.
Case 2: $\omega(a)\geq\Delta(a)$. Then $\max\{1-\frac{\omega(a)}{\Delta(a)},0\}^{2}=0$, and $N_{t}(a)>32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}$. Similar to Case 1b, we arrive at
$\text{rad}_{t}(a)=\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a)}}<\sqrt{\frac{2\log(4Kt^{4})}{32\cdot\frac{\log(4Kt^{4})}{\Delta(a)^{2}}}}\leq\frac{\Delta(a)}{4}.$
Consequently,
$\text{UCB}_{t}(a)\leq\mu^{\text{(on)}}(a)+2\cdot\frac{\Delta(a)}{4}<\mu^{(\text{on})}_{*}\leq\min\left\{\text{UCB}_{t}(a_{*}),\text{UCB}^{\text{S}}_{t}(a_{*})\right\},$
where the first and the last inequalities are by the event ${\cal E}_{t}$. Altogether, $A_{t}\neq a$ with certainty.
Altogether, all cases are covered and the Lemma is proved.
B.2.2 Proof of Theorem 4.2
Applying Lemma 4.8, we get
$\mathbb{E}[\text{Reg}_{T}(\pi,T)]$ $=\sum^{T}_{t=1}\mathbb{E}\left[(\mu_{*}^{(\text{on})}-\mu^{(\text{on})}(A_{t}))\cdot\left[\mathbf{1}({\cal E}_{t})+\mathbf{1}({\cal E}^{c}_{t})\right]\right]$
$\leq\sum^{T}_{t=1}\mathbb{E}\left[(\mu_{*}^{(\text{on})}-\mu^{(\text{on})}(A_{t}))\cdot\mathbf{1}({\cal E}_{t})\right]+\Delta_{\text{max}}\sum^{T}_{t=1}\mathbb{E}\left[\mathbf{1}({\cal E}^{c}_{t})\right],$

### v2

@@ 6  Numerical Experiments
We conduct numerical experiments to validate our findings and illustrate how $V$, $T$ and $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$ affect algorithm performance. We compare MIN-UCB with three benchmarks. The first is vanilla UCB AuerCBF02 that ignores the offline data (“PURE-UCB”). The second chooses an arm with the largest $\text{UCB}^{\text{S}}_{t}(a)$ in each round defined in (8), denoted “UCBS”. The third is MonUCB in banerjee2022artificial that incorporates the offline data under $P^{\text{(off)}}=P^{\text{(on)}}$, refered to ”MONUCB”.
In all experiments, we simulate MIN-UCB and the benchmarks over various problem instances. Offline and online rewards for each arm are independently drawn from Gaussian distributions with variance 1. We set $K=10$, and without loss of generality, let $a=1$ be the unique optimal arm. For simplicity, $T_{\text{S}}(a)=T_{\text{S}}$ for all $a$ for some integer $T_{\text{S}}>0$ and $V(a)=|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ for all $a$. The online rewawrds are set as $\mu^{\text{(on)}}(1)=1$ for the optimal arm and $\mu^{\text{(on)}}(a)=0$ otherwise, implying $\Delta(a)=1$ for all sub-optimal arms. Under each instance, each algorithm is run independently for 50 trials, and we report the sample means and standard deviations of regret.
6.1  Effect Of Discrepancy
We investigate the effect of discrepancy and $V$ on algorithm performance. We fix $T_{\text{S}}=1000$, $T=10000$, and consider two experiment groups. In the first group, “Optimistic Bias”, we set $\mu^{\text{(off)}}(1)=\mu^{\text{(on)}}(1)-v$, and $\mu^{\text{(on)}}(a)=\mu^{\text{(on)}}(a)+v$ for $a>1$, meaning that the offline data overestimate the sub-optimal arms while underestimating the optimal one. In the second group, “Pessimistic Bias”, we set $\mu^{\text{(off)}}(a)=\mu^{\text{(on)}}(a)-v$ for all $a$, meaning that the offline data consistently underestimate rewards. In both groups we vary $v\in\{0.1,0.2,\cdots,1.0\}$. The results are shown in Figure 1.
(a) Optimistic Bias
(b) Pessimistic Bias
Figure 1: Effect of Discrepancy: Both magnitude and direction of bias are important.
We analyze the Optimistic Bias group, with results shown in Figure 1(a). The regret for PURE-UCB remains constant across all values of $v$, since it ignores the offline data and relies solely on online observations. In contrast, the performance of other three policies varies with $v$, as these approaches incorporate offline information in different ways. In this setting, the offline data becomes increasingly misaligned with the online environment as the bias level $v$ increases. Formally, for each sub-optimal arm $a>1$, the discrepancy measure $\omega(a)=V(a)+\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)=2v$ for $a$. When $v\leq 0.4$, $\omega(a)<\Delta(a)$, the offline data for $a$ are informative. As shown in Figure 1(a), UCBS, MONUCB, and our MIN-UCB all outperform PURE-UCB by exploiting informative offline data. However, when $v\geq 0.5$, $\omega(a)\geq\Delta(a)$, the offline data for $a$ is misleading. Then the performance of UCBS and MONUCB, both of which rely heavily and blindly trust the offline data, deteriorates rapidly. In contrast, our MIN-UCB automatically detects the unreliability of the offline data and adapts to pure online learning, achieving regret same as the PURE-UCB. This demonstrates the robustness and adaptivity of our policy. These observations align with our instance-dependent regret bound in Section 4, which shows that for each sub-optimal $a$, the offline data are helpful only when $\omega(a)<\Delta(a)$, where the saving term $\text{Sav}_{0}(a)>0$ and the performance can be improved.
We now turn to the Pessimistic Bias group, with results shown in Figure 1(b). Unlike the Optimistic Bias group, the offline data here consistently underestimate the true rewards for all arms with the same bias $v$. Despite this, the offline data still provide informative signals for identifying the optimal arm. As shown in Figure 1(b), UCBS, MONUCB, and MIN-UCB all maintain a remarkably low and stable regret across all $v$, outperforming PURE-UCB. In this setting, the discrepancy measure satisfies $\omega(a)=V(a)+\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)=0$ for all $a$ and $v$. Consequently, the condition $\omega(a)<\Delta(a)$ always holds. Moreover, since $\omega(a)$, $T_{\text{S}}(a)$ and $\Delta(a)$ are constant across different $v$, our instance-dependent bound implies that the regret of MIN-UCB remains stable and shows a large improvement over PURE-UCB, consistent with the empirical results in Figure 1(b).
Ultimately, these observations validate our theoretical findings and illustrate the importance of both the magnitude and direction of bias in determining its role in facilitating online learning.
6.2  Effect Of $T$ and $T_{\text{S}}$
We study the effect of $T$ and $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$ on algorithm performance. We set $T_{\text{S}}=10000$, $\mu^{\text{(off)}}(1)=\mu^{\text{(on)}}(1)-v$, and $\mu^{\text{(on)}}(a)=\mu^{\text{(on)}}(a)+v$ for $a>1$ and $v\in\{0.4,0.5,0.6\}$. For each $v$, we vary $T$ from 50 to 7000. The results, shown in Figure 2, reveal a clear dependence on offline data quality. When $v$ is small ($v<0.5$), the offline data are informative, UCBS, MONUCB, and MIN-UCB outperform PURE-UCB. In contrast, when $v$ is large ($v\geq 0.5$), the offline data becomes misleading. In this case, UCBS and MONUCB, which blindly rely on the information from offline data, suffer significant performance deterioration. In constrast, MIN-UCB effectively detects uninformative offline data and defaults to pure online learning, achieving the similar performance as PURE-UCB. These findings validate our regret bounds and highlight the importance of both high-quality offline data and a robust, adaptive policy that can safely and effectively utilize such data, like our MIN-UCB.
(a) $V_{\max}=0.4$
(b) $V_{\max}=0.5$
(c) $V_{\max}=0.6$
Figure 2: Effect of $T$ and $T_{\text{S}}$: informative offline data can significantly enhance online learning.
7  Conclusion and Remarks
We propose novel algorithms to harness possibly biased offline data in the stochastic multi-armed bandit and combinatorial bandit settings. Our analysis demonstrates that our algorithms achieve the best possible regret bounds under the provision of valid bias bounds, which are shown to be necessary for out-performing existing algorithms. Through theoretical results and numerical validations, we demonstrate how various factors, such as the biases’ magnitudes, the directions of the biases, valid bias bounds, affect the usefulness of offline data in online learning. We also demonstrate the applications of our framework, such as dynamic pricing and social influence maximization. Interesting future directions include extending our framework to other online learning models, such as online Markov decision processes and contextual bandits, and investigating whether the benefits and limitations of leveraging offline data remain in other models. It is also intriguing to study our setting with other metrics on the distribution drift.
@@ Appendix A More Detailed Comparisons with Existing Works
A.1  Comparing with zhang2019warm
zhang2019warm consider a stochastic contextual $K$-armed bandit model with possibly biased offline data, which generalizes our setting of stochastic $K$-armed bandits. On one hand, their proposed algorithm ARROW-CB does not require any knowledge on the discrepancy between the offline and online data, while our proposed algorithm requires knowing an upper bound $V(a)$ to $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$ for each $a$. On the other hand, in the case of stochastic $K$-armed bandits, the regret bound of ARROW-CB (see Theorem 1 in zhang2019warm) is at least the regret bound of the explore-then-commit algorithm (see Chapter 6 in lattimore2020bandit) that ignores all offline data, which in particular does not offer improved regret bound compared to existing baselines that do not utilize offline data.
@@ A.1  Comparing with zhang2019warm
The function $\text{disc}(\lambda,I)$ (see Equation (4) in zhang2019warm) is a discrepancy measure on the distance between the offline and online models, with the property that $\text{disc}(\lambda,I)\geq 0$ for all $\lambda,I$.
@@
Finally, one additional distinction is that zhang2019warm requires $T_{\text{S}}(1)=\ldots=T_{\text{S}}(K)$, while we allow the offline sample sizes $T_{\text{S}}(1),\ldots,T_{\text{S}}(K)$ to be arbitrary non-negative integers.
@@ A.2  Comparing wtih Online Learning with Advice
[…]
@@
[…]
A.3  Comparing with chen2022data
chen2022data explores a Finite-Horizon Reinforcement Learning Model incorporating possibly biased historical data, another generalization of our stochastic multi-armed bandit model. Similar to our work, their approach ”Data-pooling Perturbed LSVI” (Algorithm 2) requires an upper bound on the discrepancy between offline and online data, defined as $\Delta$ in their Assumption 1. They propose an cross-validation method and illustrate through a case study that a substantial range of $\Delta$ values produce similar performance for their algorithm. However, their approach cannot facilitate the design of an algorithm that achieve our improved regret bounds compared to existing benchmarks that neglects the offline data.
@@ A.3  Comparing with chen2022data
However, (26) does not offer an explicit closed-form bound, and they do not provide any regret lower bound.More importantly, their implicit bound solely depends on $T_{\text{S}}^{\min}=\min_{h,s,a}T_{\text{S}}(h,s,a)$, which is equivalent to $\min_{a}T_{\text{S}}(a)$ in our model. In contrast, both our instance dependent bound (Theorem 2) and instance independent bound (Theorem 4) demonstrate how the difference in $T_{S}(a)$ among different arms affects the regret. Thus, their approach cannot deliver the tight regret bounds on and insights into our model.
Appendix B Proofs for Instance-dependent Regret Upper Bounds
B.1  Proof for Lemma 1
@@
Proposition 1 (Chernoff Inequality).
@@
Proof.
Proof of Lemma 1 To prove the Lemma, it suffices to prove that $\Pr[\xi^{\text{S}}_{t}(a)]\geq 1-\delta_{t}$. Indeed, the inequality implies $\Pr(\xi_{t}(a))\geq 1-\delta_{t}$ by setting $T_{\text{S}}(a)=0$, and then a union bound on the failure probabilities of $\xi^{\text{S}}_{t}(a),\xi_{t}(a)$ for $a\in{\cal A}$ establishes the Lemma. Now, we have
@@
$\geq$ $1-\delta_{t}/2.$ (31)
The justifications of (30, 31) are the same as (28, 29) respectively. Altogether, the Lemma is proved.
B.2  Proof of instance-dependent regret upper bound for Multi-armed Bandit
B.2.1  Proof of Theorem 2
Applying Lemma 2, we get
$\mathbb{E}[\text{Reg}_{T}(\pi,T)]$ $=\sum^{T}_{t=1}\mathbb{E}\left[(\mu_{*}^{(\text{on})}-\mu^{(\text{on})}(A_{t}))\cdot\left[\mathbf{1}(\xi_{t})+\mathbf{1}(\xi^{c}_{t})\right]\right]$
$\leq\sum^{T}_{t=1}\mathbb{E}\left[(\mu_{*}^{(\text{on})}-\mu^{(\text{on})}(A_{t}))\cdot\mathbf{1}(\xi_{t})\right]+\Delta_{\text{max}}\sum^{T}_{t=1}\mathbb{E}\left[\mathbf{1}(\xi^{c}_{t})\right],$